OpenStudy (anonymous):

Solve for b. -4b + 7 ≤ 31 b ≤ -6 b < -6 b ≥ -6 b < -8

OpenStudy (trojanpoem):

b > or = to -6

OpenStudy (trojanpoem):

the 3rd answer

Nnesha (nnesha):

first subtract 7 both side

OpenStudy (trojanpoem):

\[-4b + 7 \le 31\] \[-4b \le 24\] \[b \ge -6\]

Nnesha (nnesha):

 @undeadking1357 subtract 7 both side and let me know what you get after this